# Real-to-Sim Adaptation via High-Fidelity Simulation to Control a Wheeled-Humanoid Robot with Unknown Dynamics

### v1

@@ I Introduction
Wheel-legged robots have emerged as a potential platform to facilitate operations in a wide range of industries, including agriculture, mining, military, and search and rescue. Combining the mobility of wheels and the mobility of legs, the capability to navigate difficult environments could be greatly improved. [1, 2, 3, 4].
Controlling a wheeled-legged robot poses significant challenges due to its inherent instability and complex nonlinear dynamics. Traditional methods often simplify the system by linearizing the system dynamics around its equilibrium point, enabling the use of diverse control techniques like the Linear Quadratic Regulator (LQR) or Model Predictive Control (MPC) [1, 5, tan2018sim2]. The underlying assumption is that the system can be treated as a linear system within the equilibrium point. For example, within the equilibrium point, the inverted pendulum remains upright and stationary, meaning that the net force acting on the pendulum is zero and it does not experience any acceleration. However, when the robot engages in tasks like manipulating objects or pushing heavy loads, these changes can drastically alter the equilibrium point due to shifts in the center of mass (CoM) or contact forces, making linear approximations inadequate for effective control [6]. Recently, approaches such as deep reinforcement learning (RL) and domain randomization have emerged, showing a promising result in locomotion tasks [7, 8, 9]. Yet, they suffer from interpretability issues and cannot assure stability.
@@ II Related Work
Wheeled-Legged Robot Control: The most popular approach to control wheeled-legged robots is model-based control, which leverages a mathematical dynamics model. Some strategies use simplified models, like the Wheeled Inverted Pendulum (WIP), and linearize them around the system’s equilibrium point [3, 2]. To achieve better tracking and versatility, more complex models have been applied [1, 10, 11]. However, the success of these methods depends greatly on the precision of the modeled dynamics, a precision that is difficult to attain and may not adjust well to changes in the system or environment. As an alternative, adaptive controllers have been presented to enable the robots to adapt to new environments [12, 13]. Moreover, learning-based methods, combining deep neural networks with reinforcement learning (RL), have been suggested as an effective approach to address the nonlinear locomotion challenges of legged robots [8, 7]. With the benefits of simulation for safe and efficient training, many RL studies employ sim-to-real techniques, resulting in successfully transferring the RL policy to the real world [14, 15, 16]. Despite their innovative success, they often require a lot of manual tuning or have many difficulties in interpreting the behavior and ensuring the safety of the system.
@@ III BACKGROUND
$\begin{gathered}\bigg{(}m_{o}\!+\!m_{w}\!+\!\frac{I_{w}}{r^{2}}\bigg{)}\ddot{x}_{w}\!+\!m_{o}Ls({\theta})\dot{\theta}^{2}\!-\!m_{o}Lc({\theta})\ddot{\theta}\!=\!u\\ (m_{o}L^{2}+I_{o})\ddot{\theta}-m_{o}Lc({\theta})\ddot{x}_{w}-m_{o}gLs({\theta})\!=\!0,\end{gathered}$ (2)
@@ IV-A Real-to-Sim Adaptation via a High-Fidelity Simulation
where $\alpha$ represents the damping ratio, which delays signals, and $F_{ss}$ the static and dynamic friction model, incorporating a dead zone effect denoted by $\epsilon$. The friction model, $F_{ss}$, integrates Coulomb friction $F_{c}$, the differential static friction, and viscous damping $\sigma$, where the damping force correlates with velocity. The Stribeck velocity, $vs$, marks the shift from static to dynamic friction. Notably, the friction force, $F_{ss}$, is significant in steady states with minimal movement $v$,
Figure 3: Visualizing the Nonlinear Dynamic Function $g_{\mathbf{\bm{\zeta}}}$. The selected parameter $\mathbf{\bm{\zeta}}$ is [0.15, 0.12, 0.2, 0.02, 0.01, 0.7]. The model exhibits non-responsiveness near zero speed due to the dead zone effect, with a behavior that is more nonlinear compared to the standard viscous and Coulomb friction model. Additionally, a slight phase shift is observed, indicative of a delay.
potentially leading to instability. To mitigate this, we introduce a deadzone $\epsilon$ in Eq. (5). In our application, the model $g_{\mathbf{\bm{\zeta}}}$ is specifically applied to translation and actuator joints separately to address high friction and latency issues. Moreover, we applied the artificial high-frequency noise to acquire more realistic observed state $\mathbf{\bm{X_{S}}}$. The calculation for this high-frequency noise denoted as $w$, is presented subsequently.
@@
Figure 4: Sample Result of Data Trajectory Comparison Between a Simulation and the Real World. The last graph presents a comparison of data trajectories between simulation and reality for all methods. Our approach effectively narrows the reality gap and aligns the trajectory shape with the actual one. The residual error is attributed from coupled-dynamics, non-parametric nonlinear dynamics, and class-specific data variance.

### v2

# Toward Control of Wheeled Humanoid Robots with Unknown Payloads: Equilibrium Point Estimation via Real-to-Sim Adaptation

@@ I Introduction
Wheel-legged robots have emerged as a potential platform to facilitate operations in a wide range of industries, including agriculture, mining, military, and search and rescue. Combining the mobility of wheels and the mobility of legs, the capability to navigate difficult environments could be greatly improved. [1, 2, 3, 4, 5].
Controlling a wheeled-legged robot poses significant challenges due to its inherent instability and complex nonlinear dynamics. Traditional methods often simplify the system by linearizing the system dynamics around its equilibrium point, enabling the use of diverse control techniques like the Linear Quadratic Regulator (LQR) or Model Predictive Control (MPC) [1, 6]. The underlying assumption is that the system can be treated as a linear system within the equilibrium point. For example, within the equilibrium point, the inverted pendulum remains upright and stationary, meaning that the net force acting on the pendulum is zero and it does not experience any acceleration. However, when the robot engages in tasks like manipulating objects or pushing heavy loads, these changes can drastically alter the equilibrium point due to shifts in the center of mass (CoM) or contact forces, making linear approximations inadequate for effective control [7]. Recently, approaches such as deep reinforcement learning (RL) and domain randomization have emerged, showing a promising result in locomotion tasks [8, 9, 10]. Yet, they suffer from interpretability issues and cannot assure stability.
@@ II Related Work
Wheeled-Legged Robot Control: The most popular approach to control wheeled-legged robots is model-based control, which leverages a mathematical dynamics model. Some strategies use simplified models, like the Wheeled Inverted Pendulum (WIP), and linearize them around the system’s equilibrium point [3, 2, 5]. To achieve better tracking and versatility, more complex models have been applied [1, 11, 12]. However, the success of these methods depends greatly on the precision of the modeled dynamics, a precision that is difficult to attain and may not adjust well to changes in the system or environment. As an alternative, adaptive controllers have been presented to enable the robots to adapt to new environments [13, 14]. Moreover, learning-based methods, combining deep neural networks with reinforcement learning (RL), have been suggested as an effective approach to address the nonlinear locomotion challenges of legged robots [9, 8]. With the benefits of simulation for safe and efficient training, many RL studies employ sim-to-real techniques, resulting in successfully transferring the RL policy to the real world [15, 16, 17]. Despite their innovative success, they often require a lot of manual tuning or have many difficulties in interpreting the behavior and ensuring the safety of the system.
@@ III BACKGROUND
$\begin{gathered}\bigg{(}m_{b}\!+\!m_{w}\!+\!\frac{I_{w}}{r^{2}}\bigg{)}\ddot{x}_{w}\!+\!m_{b}Ls({\theta})\dot{\theta}^{2}\!-\!m_{b}Lc({\theta})\ddot{\theta}\!=\!u\\ (m_{b}L^{2}+I_{b})\ddot{\theta}-m_{b}Lc({\theta})\ddot{x}_{w}-m_{b}gLs({\theta})\!=\!0\end{gathered}$ (2)
@@ IV-A Real-to-Sim Adaptation via a High-Fidelity Simulation
where $\alpha$ represents the damping ratio, which delays signals, and $F_{ss}$ the static and dynamic friction model, incorporating a dead zone effect denoted by $\epsilon$. The friction model, $F_{ss}$, integrates Coulomb friction $F_{c}$, the differential static friction, and viscous damping $\sigma$, where the damping force correlates with velocity (see Fig. 3). The Stribeck velocity, $vs$, marks the shift from static to dynamic friction. Notably, the friction force, $F_{ss}$, is significant in steady states with minimal movement $v$,
Figure 3: Visualizing the Nonlinear Dynamic Function $g_{\mathbf{\bm{\zeta}}}$. The selected parameter $\mathbf{\bm{\zeta}}$ is [0.15, 0.12, 0.2, 0.02, 0.01, 0.7]. The model exhibits non-responsiveness near zero speed due to the dead zone effect, with a behavior that is more nonlinear compared to the standard viscous and Coulomb friction model (Baseline). Additionally, a slight phase shift is observed, indicative of a delay.
leading to instability. To mitigate this, we introduce a deadzone $\epsilon$ in Eq. (5). In our application, the model $g_{\mathbf{\bm{\zeta}}}$ is specifically applied to translation and actuator joints separately to address high friction and latency issues. Moreover, we applied the artificial high-frequency noise to acquire more realistic observed state $\mathbf{\bm{X_{S}}}$. The calculation for this high-frequency noise denoted as $w$, is presented subsequently.
@@
Figure 4: Sample Result of Data Trajectory Comparison Between a Simulation and the Real World. The last graph presents a comparison of data trajectories between simulation and reality for all methods. Our approach effectively narrows the reality gap and aligns the trajectory shape with the actual one. The residual error is attributed from coupled-dynamics, non-parametric nonlinear dynamics, and class-specific data variance. All data samples are collected from the WIP system using LQR control.
@@ VI-C Ablation Study
VI-D Limitation
Although our method is more sample-efficient than model-free RL, it still requires a target domain dataset for Real-to-Sim adaptation. Collecting this dataset is challenging for high-dimensional systems with unknown payloads. Developing a Real-to-Sim adaptation method that does not rely on a target dataset is an interesting future direction. Additionally, to enhance simulation fidelity, errors from non-parametric models should be addressed using techniques like Gaussian Processes and Deep Neural Networks, which are commonly employed to represent such models.